OpenStudy (anonymous):

True or false? The function f(x) = 5x is an exponential function. I think it is true just want to make sure

jimthompson5910 (jim_thompson5910):

if the function is \[\large f(x) = 5^x\] then you'd be right

OpenStudy (anonymous):

thnx hb this one The function y = log2x is a logarithmic function.

OpenStudy (anonymous):

true or false

jimthompson5910 (jim_thompson5910):

you tell me

OpenStudy (anonymous):

true

jimthompson5910 (jim_thompson5910):

yep because there's a log in there

OpenStudy (anonymous):

what about this one y = 6x is written as log6y = x

OpenStudy (anonymous):

true or false

jimthompson5910 (jim_thompson5910):

\[y = 6^x\] right?

OpenStudy (anonymous):

no y=6x

jimthompson5910 (jim_thompson5910):

so x isn't the exponent?

OpenStudy (anonymous):

correct x is not an exponent

jimthompson5910 (jim_thompson5910):

ok so y = 6x is linear so logs aren't involved at all

OpenStudy (anonymous):

so it is false

OpenStudy (anonymous):

this is wht it looks like

jimthompson5910 (jim_thompson5910):

yes it's false

OpenStudy (anonymous):

ok thnx

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

is this one true

jimthompson5910 (jim_thompson5910):

look at the definition #1 on this page http://www.mathwords.com/l/logarithm_rules.htm and you tell me

OpenStudy (anonymous):

so it would be true

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

Write 7^x = 2401 as a logarithmic function.

jimthompson5910 (jim_thompson5910):

last one and you tell me

jimthompson5910 (jim_thompson5910):

use the definition you just used on the previous problem

OpenStudy (anonymous):

logx^7 = 2401 ???

jimthompson5910 (jim_thompson5910):

no what is the base of 7^x

OpenStudy (anonymous):

wat do u mean

jimthompson5910 (jim_thompson5910):

in any exponential expression, you have a base and an exponent

jimthompson5910 (jim_thompson5910):

what is the base of 7^x

OpenStudy (anonymous):

7

jimthompson5910 (jim_thompson5910):

so that's the base of the log

OpenStudy (anonymous):

so it woul be log24017 = x

OpenStudy (anonymous):

log2401^7 = x

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

7 is the base of the log

jimthompson5910 (jim_thompson5910):

so you have it backwards

OpenStudy (anonymous):

logx2401 = 7

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

no you were closer last time

OpenStudy (anonymous):

logx^7 = 2401

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

no it's \[\large \log_{7}(2401) = x\]

OpenStudy (anonymous):

thnx

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

Write log9729 = 3 in exponential form. is it 9^3